OpenStudy (anonymous):

(9y6+3)-8y6 the 6 is an exponent,what is the simplest term?

OpenStudy (anonymous):

please help

OpenStudy (danjs):

\[(9y^6+3 )- 8y^6\] that?

OpenStudy (anonymous):

yes

OpenStudy (danjs):

Since it is just addition and subtraction, you can remove the parenthesis. Then for each exponent term with a number in front, you can combine.

OpenStudy (danjs):

\[9y^6-8y^6 +3\]

OpenStudy (danjs):

\[(9-8)*y^6+3\]

OpenStudy (danjs):

\[y^6+3\]

OpenStudy (anonymous):

it was wrong

OpenStudy (danjs):

That is the simplified form. The general rule is... \[a*x^n + b*x^n = (a+b)*x^n\]

OpenStudy (danjs):

Must have typed somehting wrong here or there, for \[(9y^6+3)-8y^6 = y^6+3\]